OpenStudy (anonymous):

For 2x – 3y = –7, when x = 4, What is the value of y?

hartnn (hartnn):

just put x=4 in that so 2(4)-3y = -7 8-3y=-7 can u solve further ?

OpenStudy (anonymous):

y = 5

OpenStudy (anonymous):

yes i can solve further, thanks:)

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

solving further @hartnn -3y = -7-8 -3y = -15 3y=15 y = 15/3 y = 5

OpenStudy (anonymous):

@Ekaansh your not aloud to give them the answer

OpenStudy (anonymous):

@SugarHighGirl i just solved it further

OpenStudy (anonymous):

Yes u did but she had to

OpenStudy (anonymous):

It's her work she asked for help not the whole answer

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

but i got the answer before he solved further

OpenStudy (anonymous):

if someone gives me the answer i don't write it down right away i solve it before anything

hartnn (hartnn):

Excellent @jeyde205 !

OpenStudy (anonymous):

Ok...But he could get in trouble if u didnt solve it before he did the answer

OpenStudy (anonymous):

@jeyde205

OpenStudy (anonymous):

how could he get in trouble? i've seen alot of people tell each other answers on here? and thanks @hartnn

OpenStudy (anonymous):

Mods go looking around to see if any one gives out answers my friend got kicked off cause she gave him the answer but he solved it

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

Yea im just lookin out forevery one

OpenStudy (anonymous):

Well bye im off my SUGAR high so now imma take a nap

OpenStudy (anonymous):

lol okay, have fun